OpenStudy (anonymous):

Help please algebra. f(x)=3x-7 and g(x)=-2x-6. Find (f o g)(4) please help show me how to do this.

OpenStudy (nurali):

f(x)=3x-7 g(x)=-2x-6 f(g(x))=3(-2x-6)-7 f(g(x))=-6x-18-7 f(g(x))=-6x-25 f(g(4))=-6(4)-25 f(g(4))=-24-25 f(g(4))=-49

OpenStudy (anonymous):

thank you can you help me with 1 more?

OpenStudy (anonymous):

f(x)=-2x+4 and (g)x=-6x-7. Find f(x)-g(x)

OpenStudy (nurali):

f(x)-g(x) -2x+4-6x+7 -8x+11